OpenStudy (maheshmeghwal9):

Evaluate :- \[\tan[\frac{1}{2}\cos^{-1}\frac{3}{5}].\]

OpenStudy (experimentx):

use half angle formula of tan

OpenStudy (experimentx):

tan x=??

OpenStudy (maheshmeghwal9):

\[\tan x= \frac{2 \tan \frac{x}{2}}{1+\tan^2{\frac{x}{4}}}.\]

Parth (parthkohli):

Why not draw a triangle? :-P

OpenStudy (maheshmeghwal9):

am I ryt?

OpenStudy (experimentx):

no ... that's not correct.

OpenStudy (maheshmeghwal9):

then where is the mistake ?

hartnn (hartnn):

you can use this : \(\huge \arccos x = 2 \arctan \frac{\sqrt{1-x^2}}{1+x}\)

OpenStudy (experimentx):

x/4

OpenStudy (maheshmeghwal9):

is it x/2

OpenStudy (experimentx):

yes

OpenStudy (maheshmeghwal9):

is this correct?

OpenStudy (maheshmeghwal9):

i think it's not Please help :)

hartnn (hartnn):

only the componendo dividendo part is incorrect

hartnn (hartnn):

it'll be -tan^2x/1 =-2/8

OpenStudy (maheshmeghwal9):

ya i also thought about that

OpenStudy (maheshmeghwal9):

thanx everyone fot their help. i gt it:)